# Estimation of relative risk, odds ratio and their logarithms with guaranteed accuracy and controlled sample size ratio

### v3

@@ 4.2 Group sampling
Thus, proceeding as in Sect. 2.2.2 and taking into account (141),
@@
where $\operatorname{E}[G]\bar{\phi}$ is given by (153). Figure 18 represents the theoretical approximation (155), as well as results from simulation. As with element sampling, the efficiency for large $A$ is slightly lower than in RR and LRR (Figs. 11 and 14), and for small $A$ there is almost no difference between simulation results and theoretical approximation even if $\phi$ is large.
@@ 5.1 Element sampling
The estimation method for the LOR, $\Psi=\log(p_{1}(1-p_{2})/(p_{2}(1-p_{1})))$, has few differences compared to that presented for OR in Sect. 4. The second stage estimates $\log(p_{1}/(1-p_{1}))$ and $\log(p_{2}/(1-p_{2}))$ separately, using for each the method described in Mendo [16, section 3], which consists of two IBS procedures with the same parameter. Specifically, to estimate $\log(p_{i}/(1-p_{i}))$, $i=1,2$, IBS is applied to population $i$ until $s_{i}$ successes are obtained, which requires a random number $N_{i}^{\prime}$ of observations from this population, and then IBS is applied until $s_{i}$ failures are obtained, which requires $N_{i}^{\prime\prime}$ additional observations. These are the same steps as for OR, but with $\alpha=0$. Then, conditioned on $s_{1}$ and $s_{2}$, which are obtained in the first stage, $H_{N_{i}^{\prime}-1}+H_{N_{i}^{\prime\prime}-1}$ is a conditionally unbiased estimator of $\log(p_{i}/(1-p_{i}))$ with conditional variance less than $1/(s_{i}-5/4)$ for any $p_{i}\in(0,1)$ [16, theorem 3]. Therefore,
@@
As with previous estimators, this choice of $r_{1}$ yields average numbers of samples close to their minimum values with respect to this parameter (results not shown). The curve $c(A,r_{1},1)=0$, plotted in Fig. 3, is almost indistinguishable from that for OR; and (160) corresponds to rounding up the ordinate values of this curve, with a minimum of $3$.
@@
Particularizing (87) for $\zeta=\Psi$, with $\partial\Psi/\partial p_{1}=1/\bar{p}_{1}$, $\partial\Psi/\partial p_{2}=-1/\bar{p}_{2}$, and then using (148), (149) and the fact that $\operatorname{Var}[\hat{\Psi}]<A$, it follows that the efficiency with element sampling is bounded by (151) (with the value of $\mu_{1}$ corresponding to LOR and with $r_{1}$ computed accordingly). Figure 19 shows the results, which are very similar to those for OR (Fig. 17).
@@ 5.2 Group sampling
The efficiency with group sampling is expressed by (155), as for OR. The results, plotted in Fig. 20, are not the same as in that case (Fig. 18) because of the differences in $\mu_{1}$ and $r_{1}$, which affect $\operatorname{E}[G]$ and thus the efficiency; but they are seen to be very similar.
@@ 6 Conclusions
As an example, consider the estimation of $p_{1}p_{2}$. A straightforward approach would be to generate samples with parameter $p_{1}p_{2}$ from samples with parameters $p_{1}$ and $p_{2}$, using a procedure analogous to the Bernoulli factory described in Sect. 4, and then apply IBS to the generated samples. However, this does not give any control on the proportion of samples from the two populations, and the average number of samples required from each population scales with the inverse of $p_{1}p_{2}$. Instead, the method presented for RR and LRR can be used: two IBS processes with parameters $s_{1}$ and $s_{2}$ are respectively applied to the two populations (second stage), which requires $N_{1}$ and $N_{2}$ samples, where $s_{1}$ and $s_{2}$ are obtained from a previous pair of IBS processes with fixed parameters $r_{1}$ and $r_{2}$ (first stage). Then,
@@ Appendix A Estimation procedure and properties of the estimators
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (76) (for LRR the explicit expression (119) can equivalently be used), or obtain it by rounding up the value from Fig. 3, with a minimum of $3$. Compute $r_{2}$ as $r_{1}+\mu_{1}-\mu_{2}$.
@@
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (145) (for LOR the explicit expression (160) can equivalently be used), or obtain it by rounding up the value from Fig. 3, with a minimum of $3$. Set $r_{2}=r_{1}$.
@@
      Take a sample from population $i$. If failure, output failure. Else, take another sample from population $i$ and output the opposite of its value.
@@
      Take a sample from population $i$. If success, output failure. Else, take another sample from population $i$ and output its value.
@@ B.1 For relative risk and log relative risk
Likewise, the distribution of $M_{2}\phi$ tends to that of a gamma random variable with location parameter $r_{2}$ and scale parameter $\sqrt{\theta}$. By the continuous mapping theorem [21, theorem 2.3], the variable $X$ defined in (21) converges in distribution to the ratio of these two gamma random variables, and therefore $Y=X/\theta$ and $1/Y$ converge in distribution to beta prime random variables with parameters $r_{2}$, $r_{1}$ and $r_{1}$, $r_{2}$ respectively [8, section 4.4]. It is easy to see that $Y^{2}$ and $1/Y^{2}$ are uniformly integrable as $\phi\rightarrow 0$, which implies [6, theorem 25.12] that the variances of $Y$ and $1/Y$ converge to those of the referred beta prime distributions. Thus, approximating $\operatorname{Var}[X]$ and $\operatorname{Var}[1/X]$ by their values for $\phi\rightarrow 0$,
@@
It will be established next that (166), (168) and the right-hand side of (172) are much smaller than (171). To this end, the following observations will be useful. From (60) and (63) it stems that $r_{1}\approx r_{2}=r_{1}+\mu_{1}-\mu_{2}$ with little approximation error. In practice, the target $A$ will typically be smaller, or much smaller, than $1$. For example, a relative RMSE of $10\%$ corresponds to $A=0.01$. On the other hand, $\xi=1$ is small compared to $1/A$, and so are the values of $r_{1}$ obtained from (81); namely $r_{1}\approx\sqrt{1/A}$, as can be seen from (82) or in Fig. 3. In the example, $A=0.01$ gives $r_{1}=9$ for RR and $10$ for LRR.
@@ B.2 For odds ratio and log odds ratio
$\operatorname{Var}\left[\Delta\right]=\operatorname{E}\left[\operatorname{Var}\left[\Delta\mid\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\Delta\mid\bar{M}_{1},\bar{M}_{2}\right]\right].$ (173)
@@
$\operatorname{E}\left[\operatorname{Var}\left[\Delta\mid\bar{M}_{1},\bar{M}_{2}\right]\right]=\operatorname{E}\left[\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right],$ (174)
@@
$\begin{split}\displaystyle\operatorname{Var}\left[\operatorname{E}\left[\Delta\mid\bar{M}_{1},\bar{M}_{2}\right]\right]&=\operatorname{Var}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]\\ &\quad+2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right].\end{split}$ (175)
@@
The first stage applies IBS to samples with parameter $\bar{p}_{i}$, using a number $\bar{M}_{i}$ of those samples, of which $r_{i}$ are successes. Each sample with parameter $\bar{p}_{i}$ is generated by the Bernoulli factory described in Sect. 4.1, taking samples with parameter $p_{i}$ as inputs. With this factory, the average number of inputs needed to produce an output is $3/2$; a success output always uses $2$ inputs, and a failure output uses either $1$ or $2$ inputs. Let $\pi_{i}$ denote the probability that $2$ inputs are used, conditioned on the output being a failure. Then $3/2=2\bar{p}_{i}+(1+\pi_{i})(1-\bar{p}_{i})$, from which
@@
The term $\operatorname{Var}\left[M_{i}\mid\bar{M}_{i}\right]$ in (179) equals the variance of a binomial random variable with parameters $\bar{M}_{i}-r_{i}$ and $\pi_{i}$. Therefore, computing $\operatorname{E}[\bar{M}_{i}]$ from (11) and substituting (180),
$\operatorname{E}\left[\operatorname{Var}\left[M_{i}\mid\bar{M}_{i}\right]\right]=\pi_{i}(1-\pi_{i})\operatorname{E}[\bar{M}_{i}-r_{i}]=\frac{(1-2\bar{p}_{i})r_{i}}{4\bar{p}_{i}(1-\bar{p}_{i})}.$ (182)
@@
$\operatorname{Var}\left[\operatorname{E}\left[M_{i}\mid\bar{M}_{i}\right]\right]=\frac{(3-4\bar{p}_{i})^{2}}{4(1-\bar{p}_{i})^{2}}\operatorname{Var}\left[\bar{M}_{i}\right]=\frac{(3-4\bar{p}_{i})^{2}r_{i}}{4\bar{p}_{i}^{2}(1-\bar{p}_{i})}.$ (183)
@@ Acknowledgments
The author would like to thank an anonymous reviewer for helpful comments, the editor-in-chief Prof. Ulrike Schneider for her efficient handling of the review process, and Dr. Ben O’Neill for advice on terminology.

### v4

@@ 4.2 Group sampling
Thus, proceeding as in Section 2.2.2 and taking into account (141),
@@
where $\operatorname{E}[G]\bar{\phi}$ is given by (153). Figure 18 represents the theoretical approximation (155), as well as results from simulation. As with element sampling, the efficiency for large $A$ is slightly lower than in RR and LRR (Figures 11 and 14), and for small $A$ there is almost no difference between simulation results and theoretical approximation even if $\phi$ is large.
@@ 5.1 Element sampling
The estimation method for the LOR, $\Psi=\log(p_{1}(1-p_{2})/(p_{2}(1-p_{1})))$, has few differences compared to that presented for OR in Section 4. The second stage estimates $\log(p_{1}/(1-p_{1}))$ and $\log(p_{2}/(1-p_{2}))$ separately, using for each the method described in Mendo [16, section 3], which consists of two IBS procedures with the same parameter. Specifically, to estimate $\log(p_{i}/(1-p_{i}))$, $i=1,2$, IBS is applied to population $i$ until $s_{i}$ successes are obtained, which requires a random number $N_{i}^{\prime}$ of observations from this population, and then IBS is applied until $s_{i}$ failures are obtained, which requires $N_{i}^{\prime\prime}$ additional observations. These are the same steps as for OR, but with $\alpha=0$. Then, conditioned on $s_{1}$ and $s_{2}$, which are obtained in the first stage, $H_{N_{i}^{\prime}-1}+H_{N_{i}^{\prime\prime}-1}$ is a conditionally unbiased estimator of $\log(p_{i}/(1-p_{i}))$ with conditional variance less than $1/(s_{i}-5/4)$ for any $p_{i}\in(0,1)$ [16, theorem 3]. Therefore,
@@
As with previous estimators, this choice of $r_{1}$ yields average numbers of samples close to their minimum values with respect to this parameter (results not shown). The curve $c(A,r_{1},1)=0$, plotted in Figure 3, is almost indistinguishable from that for OR; and (160) corresponds to rounding up the ordinate values of this curve, with a minimum of $3$.
@@
Particularizing (87) for $\zeta=\Psi$, with $\partial\Psi/\partial p_{1}=1/\bar{p}_{1}$, $\partial\Psi/\partial p_{2}=-1/\bar{p}_{2}$, and then using (148), (149) and the fact that $\operatorname{Var}[\hat{\Psi}]<A$, it follows that the efficiency with element sampling is bounded by (151) (with the value of $\mu_{1}$ corresponding to LOR and with $r_{1}$ computed accordingly). Figure 19 shows the results, which are very similar to those for OR (Figure 17).
@@ 5.2 Group sampling
The efficiency with group sampling is expressed by (155), as for OR. The results, plotted in Figure 20, are not the same as in that case (Figure 18) because of the differences in $\mu_{1}$ and $r_{1}$, which affect $\operatorname{E}[G]$ and thus the efficiency; but they are seen to be very similar.
@@ 6 Conclusions
As an example, consider the estimation of $p_{1}p_{2}$. A straightforward approach would be to generate samples with parameter $p_{1}p_{2}$ from samples with parameters $p_{1}$ and $p_{2}$, using a procedure analogous to the Bernoulli factory described in Section 4, and then apply IBS to the generated samples. However, this does not give any control on the proportion of samples from the two populations, and the average number of samples required from each population scales with the inverse of $p_{1}p_{2}$. Instead, the method presented for RR and LRR can be used: two IBS processes with parameters $s_{1}$ and $s_{2}$ are respectively applied to the two populations (second stage), which requires $N_{1}$ and $N_{2}$ samples, where $s_{1}$ and $s_{2}$ are obtained from a previous pair of IBS processes with fixed parameters $r_{1}$ and $r_{2}$ (first stage). Then,
@@ Appendix A Estimation procedure and properties of the estimators
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (76) (for LRR the explicit expression (119) can equivalently be used), or obtain it by rounding up the value from Figure 3, with a minimum of $3$. Compute $r_{2}$ as $r_{1}+\mu_{1}-\mu_{2}$.
@@
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (145) (for LOR the explicit expression (160) can equivalently be used), or obtain it by rounding up the value from Figure 3, with a minimum of $3$. Set $r_{2}=r_{1}$.
@@
      Take a sample from population $i$. If it is a failure, output failure. Else, take another sample from population $i$ and output the opposite of its value.
@@
      Take a sample from population $i$. If it is a success, output failure. Else, take another sample from population $i$ and output its value.
@@ B.1 For relative risk and log relative risk
Likewise, the distribution of $M_{2}\phi$ tends to that of a gamma random variable with location parameter $r_{2}$ and scale parameter $\sqrt{\theta}$. By the continuous mapping theorem [21, theorem 2.3], the variable $X$ defined in (21) converges in distribution to the ratio of these two gamma random variables, and therefore $Y=X/\theta$ and $1/Y$ converge in distribution to beta prime random variables with parameters $r_{2}$, $r_{1}$ and $r_{1}$, $r_{2}$ respectively [8, section 4.4]. It is easy to see that $Y^{2}$ and $1/Y^{2}$ are uniformly integrable as $\phi\rightarrow 0$, which implies [6, theorem 25.12] that the variances of $Y$ and $1/Y$ converge to those of the referred beta prime distributions. Thus, approximating $\operatorname{Var}[X]$ and $\operatorname{Var}[1/X]$ by their values for $\phi\rightarrow 0$, the following expressions are obtained:
@@
It will be established next that (166), (168) and the right-hand side of (172) are much smaller than (171). To this end, the following observations will be useful. From (60) and (63) it stems that $r_{1}\approx r_{2}=r_{1}+\mu_{1}-\mu_{2}$ with little approximation error. In practice, the target $A$ will typically be smaller, or much smaller, than $1$. For example, a relative RMSE of $10\%$ corresponds to $A=0.01$. On the other hand, $\xi=1$ is small compared to $1/A$, and so are the values of $r_{1}$ obtained from (81); namely $r_{1}\approx\sqrt{1/A}$, as can be seen from (82) or in Figure 3. In the example, $A=0.01$ gives $r_{1}=9$ for RR and $10$ for LRR.
@@ B.2 For odds ratio and log odds ratio
$\operatorname{Var}\left[\Delta\right]=\operatorname{E}\left[\operatorname{Var}\left[\Delta\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\Delta\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right].$ (173)
@@
$\operatorname{E}\left[\operatorname{Var}\left[\Delta\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]=\operatorname{E}\left[\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right],$ (174)
@@
$\begin{split}\displaystyle\operatorname{Var}\left[\operatorname{E}\left[\Delta\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]&=\operatorname{Var}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]\\ &\quad+2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right].\end{split}$ (175)
@@
The first stage applies IBS to samples with parameter $\bar{p}_{i}$, using a number $\bar{M}_{i}$ of those samples, of which $r_{i}$ are successes. Each sample with parameter $\bar{p}_{i}$ is generated by the Bernoulli factory described in Section 4.1, taking samples with parameter $p_{i}$ as inputs. With this factory, the average number of inputs needed to produce an output is $3/2$; a success output always uses $2$ inputs, and a failure output uses either $1$ or $2$ inputs. Let $\pi_{i}$ denote the probability that $2$ inputs are used, conditioned on the output being a failure. Then $3/2=2\bar{p}_{i}+(1+\pi_{i})(1-\bar{p}_{i})$, from which
@@
The term $\operatorname{Var}[M_{i}\mid\bar{M}_{i}]$ in (179) equals the variance of a binomial random variable with parameters $\bar{M}_{i}-r_{i}$ and $\pi_{i}$. Therefore, computing $\operatorname{E}[\bar{M}_{i}]$ from (11) and substituting (180),
$\operatorname{E}\left[\operatorname{Var}\left[M_{i}\mathrel{}\middle|\mathrel{}\bar{M}_{i}\right]\right]=\pi_{i}(1-\pi_{i})\operatorname{E}[\bar{M}_{i}-r_{i}]=\frac{(1-2\bar{p}_{i})r_{i}}{4\bar{p}_{i}(1-\bar{p}_{i})}.$ (182)
@@
$\operatorname{Var}\left[\operatorname{E}\left[M_{i}\mathrel{}\middle|\mathrel{}\bar{M}_{i}\right]\right]=\frac{(3-4\bar{p}_{i})^{2}}{4(1-\bar{p}_{i})^{2}}\operatorname{Var}\left[\bar{M}_{i}\right]=\frac{(3-4\bar{p}_{i})^{2}r_{i}}{4\bar{p}_{i}^{2}(1-\bar{p}_{i})}.$ (183)
@@ Acknowledgments
The author would like to thank Dr. Ben O’Neill for advice on terminology.